If a body have a mass 1000kg and moves with velocity of 34m/s. What is the momentum?

Grade:12th pass

1 Answers

Phanindra
5 years ago
Momentum , p = mv... m = mass and v = velocity... Here....m = 1000 and v = 34m/ s...So...momentum = 1000 × 34 = 34000 kg m/s....